Suppose an investor is given an investment opportunity that will provide a future value of Kshs 50,000 when it matures at the end of two years. How much would an earning of Kshs 50,000 be at the end of 2 years, at a rate of 10 percent?


1
Expert's answer
2022-02-20T16:22:53-0500

"FV=PV(1+i)^n"


"=Kshs50,000(1+0.1)^2"


"=50,000(1.1)^2= 60500"


"FV=Kshs60500"